If someone could help me with my Operations Management homework that would be great!A renovation of the gift shop at Orlando Amway Center has six activities (in hours). In the following table, the estimates of
Answer #1

PERT refers to Program evaluation and review technique where networks are used to plan and control the project and CPM refers to critical path method where shortest path is taken to plan the project. The objectives of PERT/ CPM are: -

  • Plan is made visually that help the project team to understand the progress and direction.
  • Project completion time can be determined.
  • Brainstorming what needs to be done.
  • To identify the task that require tight control.
  • To identify task that is critical for completion time.

(et 6 6 a0 2 15 00 19 2 6 32 00 18t(4x18) t18 6 18+12+18 Etc 102 18:006 = ら+52 +19 6 79 Et 17 + 11 +12 +20 1817 F 16 +(4x19+ 22 16+ 16 + 22 4 19.00 bet ui 15 00 32 00 18 00 1317 18.11 14 0 0tandard duwiatan(s)b a 6 i33 4 I 21 2-33 주 = 16-13 83 6022- 16 6 -100 ah an l 33 2.33 83 0 50 1 00